# Generative Human Geometry Distribution

### v2

@@ 4.3 Conditional Distribution Encoding
The dataset for our task consists of a set of pairs $\mathcal{D}=\{(\mathcal{S},\mathcal{T})\}$, each associated with a latent representation $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$, which is used to recover $\mathcal{T}$ by conditioning on $\mathcal{S}$. This design aligns with our generative task, i.e., $\mathcal{S}$-conditioned $\mathcal{T}$ generation. The training loss can be written as,
$\operatorname*{arg\,min}_{\theta,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}$ $\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}$ (5)
$\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathbf{z}_{\mathcal{T}|\mathcal{S}})-(\Delta\mathbf{x}-\mathbf{n})\right\|,$
To enhance the expressiveness, we model $\mathbf{z}_{\mathcal{T}|\mathcal{S}}\in\mathbb{R}^{C\times H\times W}$ as a compressed 2D feature map (height $H$ and width $W$), as shown in Fig. 11 (a). The representation aligns with recent advancements in UV representations for 3D objects [46]. Furthermore, we use a decoder network $\mathrm{Dec}_{\phi}(\cdot)$ to decompress $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$ to a higher resolution. By leveraging the UV coordinates associated with $\mathbf{x}_{0}^{\prime}$, we can then obtain per-point latents on the surface $\mathcal{S}$ by sampling the decompressed map, denoted as $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Instead of directly using $\mathrm{z}_{\mathcal{T}|\mathcal{S}}$ as input, we adopt the local geometric feature $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Thus, the revised optimization can be written as
$\operatorname*{arg\,min}_{\theta,\phi,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}$ $\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}$ (6)
$\left\|u_{\theta}\left(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})\right)-(\Delta\mathbf{x}-\mathbf{n})\right\|.$
@@
The design of the denoiser $u_{\theta}$ is adapted from [52]. In this case, we have two additional conditioning signals, SMPL point $\mathbf{x}^{\prime}_{0}$ and latent $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Additionally, we augment $\mathbf{x}^{\prime}_{0}$ by concatenating normals and canonical coordinates. Surface normals provide directional cues for clothing inference, while canonical coordinates encode body part semantics (e.g., distinguishing limbs from the torso).
@@ 4.4 Generative Tasks
When we finish the learning of the latents $\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}$, we train generative models in the latent space. We investigate two tasks: pose-conditioned generation and novel pose generation from a single-view normal image, both built on a U-Net backbone [15, 16] architecture, as illustrated in Fig. 11 (b).
Pose-conditioned generation synthesizes diverse human geometries conditioned on an SMPL template mesh. To encode pose information, we render SMPL vertex positions into UV maps and inject them as residual connections into the U-Net architecture. For novel pose generation, the model additionally takes a frontal normal image as an extra condition alongside the template mesh. Specifically, for each animation sequence in the dataset, we randomly select one frame to provide the normal image condition and another frame to provide the SMPL pose condition. We leverage the frozen DINO-ViT model [4] to extract image features, which are then fused into the U-Net via cross-attention layers.
Notably, while we explicitly condition both tasks on human pose to model the inherent clothing deformation with body articulation, our model retains the flexibility to synthesize plausible geometries even when pose information is mismatched to the feature map, as demonstrated in the “Pose Variation” case of Fig. 8.
The pose-conditioned generation networks are trained on the THuman2 dataset [49], while the novel pose generation is trained on 4DDress [33]. See the Appendix for implementation details.
@@ 5 Experiments and Results
In this section, we first conduct ablation studies on the training pair construction strategy, different formulations of geometry distribution, and network architecture. Then we compare our method with state-of-the-art generative methods regarding the pose-conditioned generation. Finally, we showcase our synthesizing results of novel pose generation.
@@ 5.1.1 Training Pairs Construction
Rather than searching for the nearest points on the dense SMPL mesh, we first sample a relatively sparse set of points from the SMPL mesh and then determine the nearest points $\mathbf{x}_{0}’$ ’ from this sampled set. It distributes the mapping workload across different $\mathbf{x}_{0}’$ ’, which improves sampling efficiency. To validate this, we train our auto-decoder under both strategies and present the results in Fig. 6.
@@
The first row shows the results using the points-set sampling strategy, while the second row corresponds to directly identifying the nearest points on the SMPL mesh. The normal maps in the second row exhibit noticeable holes, indicating insufficient sampling. These artifacts are particularly pronounced for loose clothing (middle case in Fig. 6) but also appear in regions of relatively tight clothing (left and right case in Fig. 6).
5.1.2 Geometry Distribution Formulations
Figure 4: Single-geometry fitting visualization of different distribution formulations at the $10,000$th step.
@@ 5.1.2 Geometry Distribution Formulations
To evaluate the training efficiency of different geometry distributions, we first conduct a single-geometry fitting task and then extend our analysis to a dataset. We compare our method with several alternatives, including the original geometry distribution proposed by Zhang et al. [52] (Original), the naive formulation in Eq. 2 (w/o pairs), and the formulation sampling from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$, proposed in Eq. 3.
@@
For the single-geometry fitting task, we train each formulation for $10,000$ iterations and visualize the synthesized point clouds at the final step in Fig. 4. Our method produces the smoothest and most accurate reconstructions, while all other methods generate points that deviate from the surface. The Chamfer distances at the final step are presented in Tab. 2 (Single). The original geometry distribution performs the worst because it samples from Gaussian noise and requires more iterations to capture the coarse human shape. The “w/o pairs” formulation also exhibits suboptimal performance as it must learn mappings between distant points, leading to slower convergence. The approach that samples from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ achieves the lowest Chamfer distance in this task. With a fixed pose, each sample can focus on learning specified local features, thereby simplifying the optimization. However, as shown in Fig. 4, its reliance on dispersed Gaussian centers leads to noisy output for some regions. Furthermore, this approach negatively impacts convergence when scaling to datasets with diverse poses, as reflected in its degraded performance in Tab. 2 (Dataset).
@@
For dataset-scale evaluation, we train each model on the THuman2 [49] dataset for $30,000$ iterations. As shown in Fig. 5, our method consistently achieves significantly higher fidelity compared to the other approaches within the same number of training iterations. The evolution of Chamfer distance over iterations is provided in the Appendix.
Original w/o pairs $\mathcal{N}(\mathbf{x}_{0}^{\prime},\boldsymbol{\sigma})$ Ours
@@
Table 2: Chamfer distance of different distribution formulations.
Figure 5: Visualization results of different geometry distribution formulations at the $30,000$th training iteration. GT represents the ground-truth result. Notably, the results do not represent the final outcome, as training is not fully completed.
5.1.3 Network Architecture
We experiment to validate that the auto-decoder outperforms the existing auto-encoders for our task. The auto-encoder learns features from the input, requiring more computations and being potentially constrained by the input’s representation capacity.
We adopt the 3D representation network VectorSet [50] as the encoder for the auto-encoder and take $50,000$ input points, balancing representation capability and memory consumption. Since VectorSet uses independent embeddings rather than a feature map, we treat each pixel of the feature map $\mathbb{R}^{24\times 24\times 8}$ as an individual embedding, yielding 576 embeddings (FeatureMap). To further ensure a fair comparison, we introduce an additional model where we replace our feature map decompression module (Fig. 11 (b)) with the one proposed by VectorSet, maintaining as much of the VectorSet architecture as possible (VectorSet).
The experiment is conducted on the THuman2 dataset [49]. As shown in Tab. 3, we report the average distance between the synthesized points and the surface. While converting independent embeddings to a feature map improves reconstruction accuracy, the results remain inferior to our proposed auto-decoder.
@@ 5.1.3 Network Architecture
Table 3: Surface distance comparison between various designs.
Figure 6: Comparison of strategies for constructing training pairs. The first row shows the results of our proposed strategy, and the second row finds the nearest points on the SMPL mesh.
@@ 5.2 Pose-conditioned Random Generation
This experiment measures the geometry quality of pose-conditioned generation. Consistent with previous work [57, 56], we utilize the THuman2 Dataset [49] for training and evaluation. We calculate the Fréchet Inception Distance (FID) metrics between the rendered normal images and ground truth normal maps. Notably, some works [41, 57, 56] enhance their rendering results using information such as normal field/map or predicted points’ rotations. Since our goal is to compare geometry quality, we primarily evaluate these methods based on their raw geometry outputs. To ensure fairness, we also report their performance on their final rendered results. For each subject, we render $50$ images from different views and synthesize $25,000$ normal images for the dataset.
We compare our method with representative human generation methods employing different 3D representations, including Gaussian splatting (E3Gen [56]), implicit function (GetAvatar [8], gDNA[41]), and Nerf (ENARF [24], GNARF [3] and Eva3D [11]). As shown in Tab. 4, our method significantly outperforms other approaches in terms of raw geometry outputs (blue background metric). In comparison to the state-of-the-art (gDNA), our method improves geometry quality by $57\%$ ($42.9$ to $16.2$). Additionally, when compared to the final rendered results from other methods, our raw geometry shows improvements as well ($7\%$ from $17.4$ to $16.2$). As shown in Fig. 7, E3Gen [56] synthesizes unnatural shapes with inconsistent normal colors. GetAvatar [57] generates cloth wrinkles with unnatural directional patterns, while gDNA [41]’s normals produce wrinkles that appear random, unrealistic, and not coordinated with the human pose. In contrast, our method generates realistic and pose-consistent clothing details, demonstrating superior geometric fidelity and detail preservation.
Method FID (normal) $\downarrow$
ENARF* 223.72
GNARF* 166.62
EVA3D* 60.37
E3Gen (raw geometry) 65.32
E3Gen (enhanced rendering) 28.12
GetAvatar (raw geometry) 56.07
GetAvatar (enhanced rendering) 22.77
gDNA (raw geometry) 42.90
gDNA (enhanced rendering) 17.43
Ours 16.16
Table 4: Comparison of FID scores on THuman2 dataset. The * results are adopted from E3Gen [56]. Metrics with a blue background are computed from the rendered results of the generated geometries without additional rendering enhancements.
Figure 7: Qualitative comparisons of pose-conditioned human generation results. The left two columns display final rendering results for baseline methods, while the right two columns show their raw geometry outputs. The last row illustrates two samples generated by our method.
@@ 5.3 Novel Pose Generation
Previous methods [57, 41, 56] synthesize 3D humans in canonical space and obtain novel pose appearances via deformation. As a result, while pose transformation is inherently stable, these methods lack the ability to generate pose-dependent clothing deformations (e.g., wrinkles and folds), as their geometry details remain unchanged after deformation. See the Appendix for visualization.
Unlike these methods, our approach directly synthesizes points on the deformed shape without relying on rigging mechanisms to ensure plausibility under pose changes. However, when the feature map is fixed and only the pose is altered, our method still produces reasonable results, as shown in Fig. 8 (Fixed Feature), demonstrating its robustness in handling novel poses. Furthermore, when provided with a single-view normal map specifying the clothing type, our network effectively learns a pose-aware feature map, enabling the synthesis of realistic clothing deformations, as shown in Fig. 8.
Figure 8: Comparison of novel pose generation results: (left) the source identity; (middle) novel poses generated with the source identity’s feature maps; (right) novel poses generated with the synthesized feature map from a single-view image.
In Fig. 9, we present an extreme case of novel pose generation, where the body heavily occludes most clothing, the target poses differ significantly from the input, and the clothing drapes far from the body. More results can be seen in the Appendix.
Figure 9: Extreme case of novel pose generation.
6 Limitations and Future Work
Similar to other generative methods, our approach is constrained by the diversity of the training datasets. Additionally, we cannot guarantee uniform sampling on the target geometry surface. This is because the number of target points $\mathbf{x}_{1}$ associated with each SMPL point $\mathbf{x}_{0}^{\prime}$ is not uniform. Some $\mathbf{x}_{0}^{\prime}$ may correspond to more $\mathbf{x}_{1}$ points than others. While sufficient sampling and discarding points that are too close to each other alleviate this problem, we encourage future work to explore more advanced strategies for constructing training pairs.

### v3

@@ 4.3 Conditional Distribution Encoding
The dataset for our task consists of a set of pairs $\mathcal{D}=\{(\mathcal{S},\mathcal{T})\}$, each associated with a latent representation $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$, which is used to recover $\mathcal{T}$ (target geometry) by conditioning on $\mathcal{S}$ (SMPL). This design aligns with our generative task, i.e. , $\mathcal{S}$-conditioned $\mathcal{T}$ generation. The training loss can be written as,
$\operatorname*{arg\,min}_{\theta,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathbf{z}_{\mathcal{T}|\mathcal{S}})-(\Delta\mathbf{x}-\mathbf{n})\right\|,$ (5)
To enhance the expressiveness, we model $\mathbf{z}_{\mathcal{T}|\mathcal{S}}\in\mathbb{R}^{C\times H\times W}$ as a compressed 2D feature map (height $H$ and width $W$), as shown in Fig. 4 (a). The representation aligns with recent advancements in UV representations for 3D objects Yan et al. (2024). Furthermore, we use a decoder network $\mathrm{Dec}_{\phi}(\cdot)$ to decompress $\mathbf{z}_{\mathcal{T}|\mathcal{S}}$ to a higher resolution. By leveraging the UV coordinates associated with $\mathbf{x}_{0}^{\prime}$, we can then obtain per-point latents on the surface $\mathcal{S}$ by sampling the high-resolution map, denoted as $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Thus, the revised optimization can be written as
$\operatorname*{arg\,min}_{\theta,\phi,\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}}\mathbb{E}_{(\mathcal{S},\mathcal{T})\in\mathcal{D}}\mathbb{E}_{\mathbf{n},(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1})}\left\|u_{\theta}\left(\mathbf{x}_{t},t|\mathbf{x}_{0}^{\prime},\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})\right)-(\Delta\mathbf{x}-\mathbf{n})\right\|.$ (6)
@@
The design of the denoiser $u_{\theta}$ is adapted from Zhang et al. (2025). In this case, we have two additional conditioning signals, SMPL point $\mathbf{x}^{\prime}_{0}$ and latent $\mathrm{Dec}_{\phi}(\mathbf{z}_{\mathcal{T}|\mathcal{S}})(\mathbf{x}_{0}^{\prime})$. Additionally, we augment $\mathbf{x}^{\prime}_{0}$ by concatenating normals and canonical coordinates. Surface normals provide directional cues for clothing inference, while canonical coordinates encode body part semantics (e.g., distinguishing limbs from the torso).
@@ 4.4 Generative Tasks
After finishing learning the latents $\{\mathbf{z}_{\mathcal{T}|\mathcal{S}}\}$, we train generative models in the latent space. We investigate two tasks: pose-conditioned random generation and novel pose generation of a given avatar, both built on a U-Net Karras et al. (2024; 2025) architecture, as illustrated in Fig. 4 (b).
Pose-conditioned random generation synthesizes diverse human geometries conditioned on an SMPL template mesh. To encode pose information, we render SMPL vertex positions into UV maps and inject them as residual connections into the U-Net architecture. For novel pose generation, the model additionally takes a frontal normal image as an extra condition to indicate the avatar identity. Specifically, for each animation sequence in the dataset, we randomly select one frame to provide the normal image condition and another frame to provide the SMPL pose condition. We leverage the frozen DINO-ViT model Caron et al. (2021) to extract image features, which are then fused into the U-Net via cross-attention layers.
The pose-conditioned generation networks are trained on the THuman2 dataset Yu et al. (2021b), while the novel pose generation is trained on 4DDress Wang et al. (2023). See the Appendix for implementation details.
@@ 5 Experiments and Results
In this section, we first conduct ablation studies on the training pair construction strategy and network architecture. Next, we compare different formulations of geometry distribution, including the formulation proposed by  Zhang et al. (2025) and our proposed alternatives. Finally, we compare our method with state-of-the-art generative methods regarding the two selected generative tasks.
@@ 5.1.1 Training Pairs Construction
Rather than searching for the nearest points on the dense SMPL mesh, we first sample a relatively sparse set of points from the SMPL mesh and then determine the nearest points $\mathbf{x}_{0}$’ from this sampled set. It distributes the mapping workload across different $\mathbf{x}_{0}$’, which improves sampling efficiency. To validate this, we train our auto-decoder under both strategies and present the results in Fig. 7.
@@
The first row shows the results using the sparse points set sampling strategy, while the second row corresponds to directly identifying the nearest points on the SMPL mesh. The normal maps in the second row exhibit noticeable holes, indicating insufficient sampling. These artifacts are particularly pronounced for loose clothing (middle case in Fig. 7) but also appear in regions of relatively tight clothing (left and right case in Fig. 7).
5.1.2 Network Architecture
Figure 5: Single-geometry fitting visualization.
We experiment to validate that the auto-decoder outperforms the existing auto-encoders for our task. The auto-encoder learns features from the input, requiring more computations and being potentially constrained by the input’s representation capacity. We adopt the 3D representation network VecSet Zhang et al. (2023a) as the encoder for the auto-encoder and take $50,000$ input points, balancing representation capability and memory consumption. Since VecSet uses independent embeddings rather than a feature map, we treat each pixel of the feature map $\mathbb{R}^{24\times 24\times 8}$ as an individual embedding, yielding 576 embeddings (FeatureMap). To further ensure a fair comparison, we introduce an additional model where we replace our feature map decompression module (Fig. 4 (b)) with the one proposed by VecSet, maintaining as much of the VecSet architecture as possible (VecSet).
The experiment is conducted on the THuman2 dataset Yu et al. (2021b). As shown in Tab. 3, we report the average distance between the synthesized points and the surface. While converting independent embeddings to a feature map improves reconstruction accuracy, the results remain inferior to our proposed auto-decoder.
5.2 Comparisons of Geometry Distribution Formulations
@@
To evaluate the training efficiency of different geometry distributions, we first conduct a single-geometry fitting task and then extend our analysis to a dataset. We compare our method with several alternatives, including the geometry distribution proposed by Zhang et al. (2025), the naive formulation in Eq. 2 (w/o Pairs), and the formulation sampling from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ without distribution normalization (w/o DistNorm), proposed in Eq. 3.
@@
For the single-geometry fitting task, we train each formulation for $10,000$ iterations and visualize the synthesized points at the final step in Fig. 5. Our method produces the smoothest and most accurate reconstructions, while all other methods generate points that deviate from the surface. The Chamfer distances at the final step are presented in Tab. 3 (Single). The distribution proposed by Zhang et al. (2025) performs the worst because it samples from Gaussian noise and requires more iterations to capture the coarse human shape. The “w/o Pairs” formulation also exhibits suboptimal performance as it must learn mappings between distant points, leading to slower convergence. The approach that samples from $\mathcal{N}(\mathbf{x}_{0}^{\prime},\bm{\sigma})$ (w/o DistNorm) achieves the lowest Chamfer distance in this task. With a fixed pose, each sample can focus on learning specified local features, thereby simplifying the optimization. However, as shown in Fig. 5, its reliance on dispersed Gaussian centers leads to noisy output for some regions. Furthermore, this approach negatively impacts convergence when scaling to datasets with diverse poses, as reflected in its degraded performance in Tab. 3 (Dataset).
@@
For dataset-scale evaluation, we train each model on the THuman2 Yu et al. (2021b) dataset for $30,000$ iterations. Our method consistently achieves significantly higher fidelity compared to the other approaches within the same number of training iterations, as shown in Fig. 7. Notably, the results do not represent the final outcome, as training is not fully completed. The evolution of Chamfer distance over iterations is provided in the Appendix.
Figure 6: Comparison of strategies for constructing training pairs. The first row shows our proposed strategy, the second row finds the nearest points on the SMPL mesh.
Figure 7: Visualization results of different geometry distribution formulations at the $30,000$th training iteration. GT represents the ground-truth result.
Table 2: Chamfer distance of different distribution formulations.
Zhang et al. w/o Pairs w/o DistNorm Ours
@@
Table 3: Surface distance comparison between various designs.
@@ 5.3 Pose-conditioned Random Generation
This experiment measures the geometry quality of pose-conditioned random generation. Consistent with previous work Zhang et al. (2023b; 2024d), we utilize the THuman2 Dataset Yu et al. (2021b) for training and evaluation. We calculate the Fréchet Inception Distance (FID) metrics between the rendered normal images and ground truth normal maps. Notably, some works Xu et al. (2022); Zhang et al. (2023b; 2024d) enhance their rendering results using information such as normal field/map or predicted points’ rotations. Since our goal is to compare geometry quality, we primarily evaluate these methods based on their raw geometry outputs. To ensure fairness, we also report their performance on their enhanced rendered results. For each subject, we render $50$ images from different views and synthesize $25,000$ normal images for the dataset.
We compare our method with representative human generation methods employing different 3D representations, including Gaussian splatting (E3Gen Zhang et al. (2024d)), implicit function (GetAvatar Feng et al. (2023), gDNAXu et al. (2022)), and Nerf (ENARF Noguchi et al. (2022), GNARF Bergman et al. (2022) and Eva3D Hong et al. (2022)). As shown in Tab. 4, our method significantly outperforms other approaches in terms of raw geometry outputs. In comparison to the state-of-the-art (gDNA), our method improves geometry quality by $57\%$ ($42.9$ to $16.2$). Additionally, when compared to the enhanced rendered results from other methods, our raw geometry shows improvements as well ($7\%$ from $17.4$ to $16.2$). As shown in Fig. 8, E3Gen Zhang et al. (2024d) synthesizes unnatural shapes with inconsistent normal colors. GetAvatar Zhang et al. (2023b) generates cloth wrinkles with unnatural directional patterns, while gDNA Xu et al. (2022)’s normals produce wrinkles that appear random, unrealistic, and not coordinated with the human pose. In contrast, our method generates realistic and pose-consistent clothing details, demonstrating superior geometric fidelity and detail preservation.
Table 4: Comparison of FID scores. The * results are adopted from E3Gen Zhang et al. (2024d). For some methods, the results are rendered directly from their raw geometries, so the numbers in both rows are identical.
ENARF* GNARF* EVA3D* E3Gen GetAvatar gDNA Ours
Raw Geometry 223.72 166.62 60.37 65.32 56.07 42.90 16.16
Enhanced Rendering 223.72 166.62 60.37 28.12 22.77 17.43 16.16
Figure 8: Qualitative comparisons of human generation results. Existing methods are displayed with both rendering results and raw geometry outputs, typically organized as two columns each. We include two geometry samples generated by our method for comparison.
@@ 5.4 Novel Pose Generation
Table 5: User study on quality and physical plausibility.
Metric GetAvatar gDNA E3Gen Ours
Quality ↑ 2.16 2.54 2.12 4.04
Plausibility ↑ 2.20 2.66 2.08 4.36
Previous methods Zhang et al. (2023b); Xu et al. (2022); Zhang et al. (2024d) synthesize 3D humans in canonical space and deform them via rigging, which limits pose-dependent clothing details (e.g., wrinkles and folds). In contrast, our approach directly synthesizes points on the deformed human body given the pose-aware feature map, enabling the generation of pose-dependent clothing details, as shown in Fig. 9. More visual results can be found in the Appendix.
For quantitative comparisons, FID metrics reflecting geometric quality are provided in Tab. 4. Note that since prior methods’ geometric details remain static across poses, their FID results correspond closely to those reported in the above table. In terms of physical plausibility of geometric details, due to the absence of standard metrics, we conduct a user study to demonstrate that our method uniquely enables pose-aware deformations that appear physically reasonable. For each method, we generate 2 identities under 8 poses and ask 25 participants to rate geometry quality and physical plausibility (1–5 scale), as shown in Tab. 5.
Our superior physical plausibility mainly stems from our synthesized pose-aware feature maps. Nevertheless, even when using a feature map that is incompatible with the body pose, our method still generates visually reasonable results, which demonstrates the robustness of our model in handling novel poses (see the Appendix for details).
Figure 9: Generating different poses for a given avatar. We showcase challenging cases, including exaggerated poses, skirts, and loose outfits.
6 Limitations
A limitation of our method lies in non-uniform sampling on the target geometry surface, as the number of target points $\mathbf{x}_{1}$ associated with each SMPL point $\mathbf{x}_{0}$’ varies. Although sufficient sampling and removing points that are too close mitigate this issue, we encourage future work to develop more advanced strategies for training pair construction. Besides, similar to other generative methods, our approach is constrained by the diversity of the training datasets. Specifically, our model generalizes well to various body types, as the dataset includes a reasonable diversity in this aspect. However, it cannot generate clothing styles that are entirely absent from the training data. Furthermore, the use of UV maps can cause seam artifacts in some randomly generated results due to discontinuity segmentation. A better approach would use UV segmentation aligned with real-world garment patches. Since our focus is on a modeling representation, addressing this issue is left for future work.